OpenStudy (acacia):

Solve 4x + y + z = 1 8x – 4y – 7z = 2 5y + 9z = 3 Help?

OpenStudy (luigi0210):

Matrix!

OpenStudy (luigi0210):

Do you know how to set this up as a matrix..?

OpenStudy (anonymous):

luigi!!!!

OpenStudy (luigi0210):

Yes?

OpenStudy (acacia):

I don't even have a clue @Luigi0210

OpenStudy (luigi0210):

Matrix form is the same as equation form.. just without the variables.

OpenStudy (mathstudent55):

I'd use the first two equations to eliminate x. Then use the new equation with only y and z with the third equation to solve for y and z.

OpenStudy (anonymous):

math is my dumbest point so i had seeing numbers put together like i broke my keyboard kuz i hate numbers

OpenStudy (luigi0210):

Wait, do you know how to use matrices yet? @acacia

OpenStudy (acacia):

Multipy both sides of the first equation by 2, then subract the left side of equation 2 from its left side, and E2's right side from E1's right side: 8x + 2y + 2z - 8x - (-4y) - (-7z) = 2 - 2 6y + 9z = 0 6y = -9z y = -1.5z Substitute that for "y" in E3: 5 * (-1.5z) - 9z = 3 -7.5z - 9z = 3 -16.5z = 3 z = –3/16.5 = –6/33 = –2/11 And "y" comes from y = –1.5z = (–1.5)(–2/11) = 3/11 And put that back into E2: 8x - 4(3/11) - 7(–2/11) =2 8x - 12/11 + 14/11 = 22/11 8x + 2/11 = 22/11 8x = 20/11 x = 20/88 = 5/22 This is how my answer is turning out and I don't know if its right, it looks wrong to me

OpenStudy (acacia):

@Luigi0210 sorta, kinda I'm a slow child

OpenStudy (mathstudent55):

4x + y + z = 1 8x – 4y – 7z = 2 5y + 9z = 3 Multiply the first equation by -2 Write the second equation below it Then add the two equations -8x - 2y - 2z = -2 8x - 4y - 7z = 2 ------------------- -6y - 9z = 0 Now you have a new equation in only y and z. Solve it together with he third equation, which only has y and z. -6y - 9z = 0 5y + 9z = 3 @acacia Can you handle a system of 2 equations in 2 unknowns?

OpenStudy (mathstudent55):

Hint: Add the two equations together. That will eliminate z, so you can find y.

OpenStudy (acacia):

so then would it be y=-3? @mathstudent55

OpenStudy (mathstudent55):

-6y - 9z = 0 5y + 9z = 3 ------------ -y = 3 y = -3 Correct!

OpenStudy (acacia):

I feel like I just made a break through!(:

OpenStudy (acacia):

@mathstudent55 can you help me with another one?

OpenStudy (acacia):

That is what I was about to ask would I have to do it the same way I solved for y?

OpenStudy (mathstudent55):

Now that you have y, use an equation that has only y and one more variable to find the other variable. Use equation 3: 5y + 9z = 3 Since y = -3, you get: 5(-3) + 9z = 3 Solve for z.

OpenStudy (mathstudent55):

Once you also have z, then you can use the first equation. Plug in y and z, and solve for x.

OpenStudy (acacia):

okay so z=2 right? @mathstudent55

OpenStudy (acacia):

and the the last one would be 0.5? If i'm right

OpenStudy (mathstudent55):

Correct. Good work.

OpenStudy (acacia):

but wait for my multiple choice answers it has A. (0.5, -3, 2) B. (1, 2, 3) C. (-3, 0.5, 2) D. (0.5, 3, 2) aren't A and C the same?